OpenStudy (anonymous):

Solve for t. (100-t)^-4 I know it is simple but I completely blanked out on how to solve these. Dx

OpenStudy (freckles):

I'm not sure what you mean by solve since then isn't an equation.

OpenStudy (freckles):

You can solve equations. But there is no equation is what I'm saying.

OpenStudy (anonymous):

sorry typo!!

OpenStudy (anonymous):

I will it from the start since it deals with "weird" decimals. It should read: .1=.4-(3.9x10^-7)(100-t)^-5 **just realized I wrote -4 instead of -5...its been one of those days

OpenStudy (freckles):

\[0.1=0.4-(3.9 \cdot 10^{-7})(100-t)^{-5}\]

OpenStudy (freckles):

first thing i would do is subtract 0.4 on both sides

OpenStudy (anonymous):

yes! and we gotta solve for t

OpenStudy (freckles):

\[0.1-0.4=-(3.9 \cdot 10^{-7})(100-t)^{-5}\]

OpenStudy (freckles):

After finding the difference of 0.1 and 0.4 I would then divide both sides by -(3.9*10^(-7))

OpenStudy (freckles):

do that for me and let me know what you have

OpenStudy (anonymous):

**769230.77=(100-t)^-5

OpenStudy (freckles):

ok

OpenStudy (freckles):

\[0.1-0.4=-0.3\\-\frac{-0.3}{3.9 \cdot 10^{-7}}=\frac{3}{39} 10^{7}=\frac{1}{13}10^7 \\ \frac{13}{10^{7}}=(100-t)^5\]

OpenStudy (freckles):

so raise both sides to the 1/5th power

OpenStudy (freckles):

this is so we cancel that exponent 5 thing

OpenStudy (anonymous):

okay got that but how do we get rid of that 5 exponent?

OpenStudy (freckles):

1/5*5=1

OpenStudy (anonymous):

oh!

OpenStudy (freckles):

\[(\frac{13}{10^{7}})^\frac{1}{5}=(100-t)^{5\cdot \frac{1}{5}}\]

OpenStudy (anonymous):

so t=99.934?

OpenStudy (freckles):

looks good \[t=100-(\frac{13}{10^7})^\frac{1}{5}\]

OpenStudy (freckles):

what an ugly answer

OpenStudy (anonymous):

right? its for a mixing problem so it makes sense lol

OpenStudy (freckles):

i guess there are no pretty answers in the real world

OpenStudy (anonymous):

you answer like all of my questions. I am so grateful!! lol

OpenStudy (freckles):

its np

OpenStudy (anonymous):

are you a mathematician or a prof???

OpenStudy (freckles):

i have taught a little